OpenStudy (astrophysics):

Been a long while since I've done this, could use some help haha. Find the mass of a wire.. @ganeshie8 @Kainui @zepdrix

OpenStudy (astrophysics):



OpenStudy (astrophysics):

From what I recall since mass is density*volume the integral should be \[m = \int\limits \rho ds \]

OpenStudy (astrophysics):

\[ds = \sqrt{\left( \frac{ dx }{ dt } \right)^2+\left( \frac{ dy }{ dt } \right)^2+\left( \frac{ dz }{ dt } \right)^2}dt\]

OpenStudy (astrophysics):

\[ds = 3\sqrt{1+4t^2+4t^4}dt\] mhm this is what I get for ds heh

OpenStudy (astrophysics):

Ah that works out \[ds = 3 \sqrt{(2t^2+1)^2}\]

OpenStudy (astrophysics):

Then our parameterization is just x = 3t so we have \[\rho = 1+t \implies m = 3\int\limits\limits_{0}^{1}(1+t)(2t^2+1)dt \] I think this is it?

OpenStudy (astrophysics):

I got m=8 if anyone cares lol, but I just want to know if the set up is correct haha.

OpenStudy (phi):

yes, that is what I got. Hopefully, that means great minds think alike.

OpenStudy (kainui):

Yeah I just worked it out and that's what I got, did the same steps as you too for the most part.

OpenStudy (astrophysics):

Haha thanks fellas!